Oxford Materials IT Committee Meeting

A meeting of the Department of Materials IT Committee was held at 2pm on Tuesday 28th Oct 2008 in the Hume-Rothery Meeting Room.


  • 1 Apologies for absence and Membership of the Committee
  • 2 Matters arising from the Minutes of the Meeting of the last meeting
  • 3 IT/AV – Matters for Discussion (attached)
  • 4 IT/AV – Head of IT Report (attached)
  • 5 IT/AV expenditure - matters for report (PJW)
  • 6 AOB
  • 7 Date and time of next meeting: Tuesday 3rd February 2008 at 2pm


See minutes below.


  • Review and update of IT Strategy
  • Review and update of IT Support Policy
  • Policy concerning Department Firewall
  • Policy concerning Vista and Office 2007
  • Update concerning generic accounts on library and print-room computers
  • Update on Groupware project


The report below summarises activities since the last meeting of ITC, as categorised by topic. The report is intended for record rather than for discussion at ITC meeting although queries may arise.

Admin Network
Generally there have been few changes. Materials0 was inaccessible from over the weekend of 17th-18th May due to a double disk failure and from 10.35am on Thursday 15th October until 8am the next morning due to a problem with disk partition addressing.

Teaching Laboratory
The microscope pc’s and the S530 pc were migrated from Materials2 to Materials0. A spare D600 laptop has been configured as a reserve laptop for the Teaching Laboratory but is kept by IT staff.

Lecture theatres / meeting rooms
No changes to lecture rooms. It is noted that failure of Materials0 causes the lecture room pc’s to become unusable.

Digital Printroom
The digital darkroom was renamed the Digital Print-room. Chris Akinola now provides day-to-day support for users of the Print-room. The desktop pc’s still have generic user accounts but it is intended that these PC’s will no-longer have generic user accounts thus each user will need to login.

The desktop pc’s still have generic user accounts but it is intended that these PC’s will no-longer have generic user accounts thus each user will need to login.

Major disruption to the MML was caused by repeated intermittent problems with the RAID file storage. Data was successfully recovered three times and the server was taken out of commission for testing. The MML is still running on temporary hardware but is about to be migrated back to the original hardware after replacing half the disks, the cables and the operating system. The new modelling chair and lecturer are due to arrive in January and November respectively.

Chao Lin worked closely with Alex Cock and Khalid Schofield over two separate projects : 1) commissioning of the ESTEEM server and storage; 2) Rewiring of the Jeol2200MCO room. The plans for reorganising the EM IT room have still not progressed.

Planned developments of the database information intended for presenting research projects, new projects available, publications, research highlights, department news, a department diary, private content for internal view only etc. are still under development.

Chris Akinola took over publications support. This year’s documentation production in September is summarised in the table below totalling almost 38,000 printed sides done using the print-room HPCLJ5550dn printer with estimated cost of 1p/side including paper costs plus several weeks of time doing collation and binding.

Title of publicationNo PagesNo Copies
MS/MEM Course Prelims Handbook7165
MS/MEM Course Finals Handbook 9765
MS/MEM Prelims lecture course synopses4065
MS/MEM Finals Core lecture synopses48 65
MS/MEM Finals Option lecture synopses3865
Part II Course Handbook4365+25
Graduate Course Handbook7865
Postgraduate lecture and training course synopses9865
Departmental Facilities and Procedures 34100

Researches in Progress document has been started and is scheduled to be completed by the end of November.

The new design for departmental headed notepaper incorporating the new Oxford brand was drafted past DC several times then became stuck over the summer awaiting approval which it has now received. An order will be placed soon. Meanwhile stocks of compliments slips were printed in-house to satisfy immediate requirements.

Several problems with networking at Begbroke were resolved by removing faulty VOIP hardware. Several switch upgrades were carried out over the summer, but several still remain. The main department backbone connection was upgraded to gigabit (£3k cost covered by £1k dept, £1k AIK, £1k MML) and the department firewall was replaced and also upgraded to gigabit.

Several hardware repair requests are outstanding and several more phones are needed. PJW presented a review of VOIP phones to DC with a request for funding to migrate the VOIP phones to the new Cisco VOIP system offered and managed by Telecoms/OUCS. It is intended to start rolling this new system into Begbroke before Christmas.

Several instances of p2p software have been detected and warnings issued. Several users received help removing viruses. Several systems were infected by malware sufficiently nasty to require reinstallation. There were several instances of problems with physical security and considerable time was spent analysing logs and checking the door entry systems.

Materials2 server has now been decommissioned and all users migrated to Materials0.
www-int has been decommissioned and services migrated to the replacement server int-webs.
The management servers oums-nomad and oums-wsus have been stable, but are next to be migrated to new hardware and the intention is to virtualize these servers.
The Emgroup1 raid array only has 10 users under AIK and still needs to be decommissioned .
Materials0 itself has had two problems affecting service availability. Two additional servers have been purchased and are being developed to take over from Materials0 later this year.

No significant changes to software. AJW and SGR are increasing usage of ABAQUS this year therefore needing 2 more license seats.

Waste Disposal
The storage of equipment prior to testing and disposal currently occupies about 80% of the IT room in 21 Banbury Road. Disposal needs to happen before Christmas. The IT stock storage has been moved to the new IT area in ETB.

IT Support Area
Chris Akinola started as the new IT support officer and is located in ETB room 10.14. The former photographic areas was rewired and redecorated (mostly after the arrival of Chris Akinola). Over the summer Paul Warren and Alex Cock gradually moved out of 154.30.06 and are now all IT staff are co-located on the ground floor of ETB and now provide a central IT support area.
The Digital Darkroom was renamed to be the Digital Print-room given the end of wet photographic services. The wet chemical equipment was disposed to Brian Gilmour being the last active user.


A meeting of the Department of Materials IT Committee was held at 2pm on Tuesday 28th Oct 2008 in the Hume-Rothery Meeting Room.
The following people attended: APD, PJW, AMC, CA, JDM, SCB, RT.

1. Apologies for Absence and Membership of the Committee

Apologies were received from Khalid Schofield who was on paternity leave. The committee congratulated him on the birth of Edward Schofield.

The committee welcomed John Murphy to the committee.

2. Matters arising from the minutes

a) Membership of the committee (Minute 1)
PJW made several recommendations to the Head of Department for a replacement for Jason Smith. John Murphy was nominated to join the committee.

b) IT disaster recovery plans (Minute 2a)
As reported below there have been several “disasters” with departmental servers since the last meeting which were successfully recovered showing that disaster recovery systems are in place and are functional.

There is still a need for general disaster recovery documentation to ensure that there is adequate provision and there is a clear structure for achieving efficient disaster recovery and ensuring business continuity.

Action: PJW

c) Website (Minute 2b)
Although a full review of research group websites has still not been undertaken, there has been considerable activity in this area over the past year. Six sites now use the CMS Made Simple system used on the department website.

Websites now hosted on the departmental web server include:

external.materials.ox.ac.ukHTML/PHPWarrenDue to merge into www.materials
fellows.materials.ox.ac.ukDrupalLinFellows handbook wiki
fighting.materials.ox.ac.ukJoomlaHaighExternal to department
fusion.materials.ox.ac.ukHTMLJohnsonSGR project
gallery.materials.ox.ac.ukPHPSchofieldPhotograph store
nots.materials.ox.ac.ukHTMLPantanayDefunct and inactive
outreach.materials.ox.ac.ukPHPHuttonSchools liaison live site
outreachpreview.materials.ox.ac.ukPHPHuttonSchools liaison development
qipirc.materials.ox.ac.ukCmsMadeSimpleIwasiewiczDevelopment Nanoqubit
rtp2009.materials.ox.ac.ukCmsMadeSimpleNichollsConference website
sophos-ext1.materials.ox.ac.ukHTMLCockSophos deployment
sophos-ext2.materials.ox.ac.ukHTMLCockSophos deployment
tmcn.materials.ox.ac.ukCmsMadeSimpleNicolosiConference website
wiki-eia.materials.ox.ac.ukDrupalLinAIK group
www-alloys.materials.ox.ac.ukPHPDrautzDGP project
www-coatings.materials.ox.ac.ukHTMLDrautzDGP project
www-em.materials.ox.ac.ukPHP/MYSQLMoldovanEM group
www.materials.ox.ac.ukCmsMadeSimpleWarrenDepartment website
www.materials.ox.ac.uk/fimCmsMadeSimpleWarrenFIM group
www.materials.ox.ac.uk/htscCmsMadeSimpleWarrenHTSC group
www-sgrgroup.materials.ox.ac.ukHTMLJohnsonSGR group
www-solar.materials.ox.ac.ukCmsMadeSimpleWattARRW group
www-stemgroup.materials.ox.ac.ukCmsMadeSimpleNellistPDN group

Other group websites hosted OUCS web servers (plain text/html – no php) include:

  • Materials Archaeology Group http://user.ox.ac.uk/~salter
  • Materials Modelling Laboratory http://user.ox.ac.uk/~mml
  • Micromechanics Group http://user.ox.ac.uk/~ajw
  • Processing of Advanced Materials http://user.ox.ac.uk/~pgrant
  • Scanning Tunnelling Microscopy Group http://user.ox.ac.uk/~stm
  • Semiconductor Group http://www-seminconductor.materials.ox.ac.uk

PJW needs to circulate questionnaire to all senior staff and fellows concerning website provision.

Action: PJW

d) Website (Minute 2d)
AOT and Kay Davies have updated the postgraduate admissions web pages and in particular the structure of the long page listing of projects available. The project descriptions themselves are still due to be served from a database in order to provide greater flexibility over presenting the same data on personal homepages.

e) Archive procedures. (Minute 2j)
IT staff have solved recent archiving requests by providing live storage. However no clear procedures exist for archiving data off departmental servers when people have left or for archiving collaborative storage areas.

PJW needs to circulate a questionnaire to all senior staff concerning archiving provision in order to find out whether there is a problem and whether we should investigate a central solution to what is currently a group-centric problem.

Action: PJW

f) Begbroke Annex network upgrade. (Minute 6b)
KS upgraded the switch in the Begbroke Highbay Annex to 100Mbit as part of network changes scheduled during the summer.

3 Matters for Discussion

a) The committee quickly reviewed the Terms of Reference of the committee as published at http://www.materials.ox.ac.uk/it/itcommittee.html and no changes were deemed necessary.

b) Committee members were requested to comment on the IT strategy and IT Support Policy at www.materials.ox.ac.uk/it/itstrategy.html www.materials.ox.ac.uk/it/itsupportpolicy.html and send any comments directly to PJW who will update them soon.

Action: PJW/All

c) Department Firewall:
Over the summer the backbone connection was upgraded to gigabit and the department firewall was replaced. PJW reported that since the firewall is mission critical the old firewall is being retained to act as a backup in the event of a hardware failure, meanwhile new firewall hardware has been ordered to configure the next generation firewall with load balancing.

The firewall upgrade caused us to review the firewall rules and exceptions. There was some discussion about existing exceptions. It was agreed that the department firewall policy should not allow blanket exceptions. It was agreed that PJW and APD should discuss this with Nicole Grobert. It was agreed that a webpage should be created in order to advertise the department firewall policy and rule sets.

Action: PJW/All

d) Policy on Vista
Proposed statements on policy about support of Vista operating systems were presented and discussed. There are still several core functions which are not properly functional under Vista. It was agreed that:

  • department machines should remain using Windows XP unless a clear justification for using Vista (only Business/Enterprise/Ultimate) is put forward by the user and accepted by IT staff
  • personally owned machines can use Vista but users should be notified that support from IT staff for problems with Vista will be low priority.

e) Policy on Office2007
Proposed statements on policy about support of Office2007 were presented and discussed. It was agreed that:

  • department IT staff should be a position to offer Office2007 upon request by Christmas (ie. should have sorted out installation, and done training for providing support for Office 2007 which is a significantly different user interface to 2003)

f) Update on Groupware Project
PJW updated the committee on the progress of the University Groupware project as detailed at http://www.ict.ox.ac.uk/odit/projects. APD requested that IT staff continue to monitor progress of the project but also proactively assess the implications and potential uses for this department. IT should also keep an eye on the Core User Directory project.

Action: PJW

g) Update on generic accounts
At the last meeting it was proposed to stop generic accounts in the library and print room. This has not been done yet but it is still an aim due to needing an audit trail.

4. Head of IT Report

A written report was circulated with the agenda. (see above)

5. IT/AV Expenditure

PJW was not ready to present financial update but will circulate details later.

6. AOB

a) Lecture Theatre and Meeting Rooms
SCB reported some pens had not been replenished in some meeting rooms. PJW to email staff and fellows to remind them to record problems in the logbook and notify itsupport@materials.ox.ac.uk.

Action: PJW

b) Oracle printed invoices SCB asked whether there was a way of getting PDF purchase orders out of oracle suitable for emailing to company contacts, other than by re-scanning the printed purchase orders. IT staff noted that Oracle prints direct to the printers not via the desktop client. AMC to investigate.

Action: AMC

c) Podcasting
SCB asked whether there was existing recordings of lectures/meetings suitable for creating a Materials area on the Oxford iTunes site. PJW reported that recordings do exist. PJW and SCB to investigate (note time limited) and talk to Schools Liaison team. Note permissions will be required and the Head of Department is required to authorise content.

Action: PJW/SCB

d) WiFi at Begbroke
APD mentioned there have been requests for networking for people visiting Begbroke (internal and external). PJW reported that when the switch roll at Begbroke is complete we will be able to implement wireless hotspots at Begbroke (EM tea area, Faraday area, APL). PJW noted that the write-up pc was decommissioned over the summer. JDM was not aware that provision had existed. PJW and KS to implement wireless and create hotdesk area(s) with access to printing facilities.

Action: PJW/SCB

e) Central File Storage
JDM noted that procedures within groups for accessing shared storage areas for files and documents are very variable. AMC noted that Materials0 is the preferred solution for shared file storage. IT staff to promote usage of Materials0 by discussing requirements with individual groups.

Action: PJW/AMC

AMC noted that IT support should try to be more service-centric and the IT web pages should be updated to more clearly promote the services available and give documentation upon how to achieve common tasks.

Action: PJW/AMC

7. Date of Next Meeting

The next meeting will take place on Tuesday 3rd February 2008 at 2pm. in Hume-Rothery Meeting Room (room booked already)

The meeting closed at 3.45pm